OpenStudy (anonymous):

6+radical 2 State the conjugate and then multiply them together. Please hurry.

jimthompson5910 (jim_thompson5910):

Hint: The conjugate of \(\large a+\sqrt{b}\) is \(\large a-\sqrt{b}\)

jimthompson5910 (jim_thompson5910):

when you multiply them together, you get \[\large (a+\sqrt{b})(a-\sqrt{b}) = a^2 - (\sqrt{b})^2\] \[\large (a+\sqrt{b})(a-\sqrt{b}) = a^2 - b\]

OpenStudy (anonymous):

So would the final answer be 6^2-(radical2)^2?

jimthompson5910 (jim_thompson5910):

you can simplify further

OpenStudy (anonymous):

6-(radical 2)?

jimthompson5910 (jim_thompson5910):

6^2 = ??

OpenStudy (anonymous):

36-(radical 2)^2?

OpenStudy (anonymous):

is it 36-(radical 2)^2

jimthompson5910 (jim_thompson5910):

The conjugate of \(\large 6+\sqrt{2}\) is \(\large 6-\sqrt{2}\) When you multiply them together, you get \[\large (6+\sqrt{2})(6-\sqrt{2}) = 6^2 - (\sqrt{2})^2\] \[\large (6+\sqrt{2})(6-\sqrt{2}) = 36 - 2\] \[\large (6+\sqrt{2})(6-\sqrt{2}) = 34\] So, Conjugate: \(\large 6-\sqrt{2}\) Answer you get when you multiply original expression by conjugate: 34

OpenStudy (anonymous):

Thanks a million onbe more question though

OpenStudy (anonymous):

(radical 2x-8)+10=6

OpenStudy (anonymous):

i needs to solve and i only have 3 minutes

jimthompson5910 (jim_thompson5910):

oh this is timed?

OpenStudy (anonymous):

i have to go and i wanna finish this

jimthompson5910 (jim_thompson5910):

well sorry I can't help with tests

OpenStudy (anonymous):

its homework, im leaving for a roadtrip and dont want to have work when i get back on sunday

jimthompson5910 (jim_thompson5910):

ok, but 3 min isn't a lot of time, sorry esp for something this complicated

jimthompson5910 (jim_thompson5910):

so maybe it's best to save it til you get back